# Zhuoming Chen Ranajoy Sadhukhan Zihao Ye Yang Zhou Jianyu Zhang Niklas Nolte

### v2

@@ Abstract
Large language models (LLMs) with long context windows have gained significant attention. However, the KV cache, stored to avoid re-computation, becomes a bottleneck. Various dynamic sparse or TopK-based attention approximation methods have been proposed to leverage the common insight that attention is sparse. In this paper, we first show that TopK attention itself suffers from quality degradation in certain downstream tasks because attention is not always as sparse as expected. Rather than selecting the keys and values with the highest attention scores, sampling with theoretical guarantees can provide a better estimation for attention output. To make the sampling-based approximation practical in LLM generation, we propose MagicPIG, a heterogeneous system based on Locality Sensitive Hashing (LSH). MagicPIG significantly reduces the workload of attention computation while preserving high accuracy for diverse tasks. MagicPIG stores the LSH hash tables and runs the attention computation on the CPU, which allows it to serve longer contexts and larger batch sizes with high approximation accuracy. MagicPIG can improve decoding throughput by 1.93.9×1.9$\sim$ 3.9$\times$1.9 ∼ 3.9 × across various GPU hardware and achieve 110ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct model with a context of 96k tokens. The code is available at https://github.com/Infini-AI-Lab/MagicPIG.
@@ 1 Introduction
Large language models (LLMs) with long context windows, such as GPT [1], Llama [25], and Gemini [59], have gained significant attention for their ability to enhance applications like chatbots [18], search engines [61], and video analysis [17]. However, serving long-context LLMs is highly challenging due to the unique bottleneck in auto-regressive generation—the key-value (KV) cache, which stores intermediate attention keys and values to avoid re-computation [50, 67]. Specifically, the KV cache grows linearly with both the batch size and sequence length, occupying substantial GPU memory and increasing decoding time. Moreover, the KV cache makes LLM generation extremely memory-bound, leading to underutilization of GPU computational power. For instance, an NVIDIA A100-40GB GPU can only handle a single request for Llama with a 128k context length, with nearly half of the decoding time spent accessing the KV cache, and poor GPU utilization  [27].
Leveraging the common insight that attention is naturally sparse, dynamic sparse or $\mathrm{TopK}$-based approximation has been extensively studied [58, 55, 66, 62], but three major challenges prevent a wide adoption in LLM serving systems. (1) Quality Degradation. They usually propose various strategies to approximate a subset of KV cache that yields the highest attention scores. However, $\mathrm{TopK}$ attention itself is a biased attention approximation and lacks theoretical guarantees. Figure 1 shows that even exact $\mathrm{TopK}$ attention results significantly degrade the accuracy of certain downstream tasks. (2) High Overhead. There is a large overhead to identify $\mathrm{TopK}$ attention, which becomes the bottleneck rather than the attention computation. For example, as studied in Liu et al. [42], naively applying a search algorithm like IVF [24] requires access over $30\%$ key states to obtain the exact $\mathrm{TopK}$, showing an unsatisfying trade-off between search accuracy and cost. (3) No Memory Saving. Although saving KV cache loading time, they cannot reduce the total memory occupied by the KV cache, which limits the maximum context and batch sizes when VRAM is scarce.
@@
  • Seemingly high sparsity is usually a consequence of an attention sink. Most of the attention scores concentrate on initial tokens (attention sink phenomenon) [63], making the distribution look sparser. However, as shown in Figure 2b, attention scores are distributed more uniformly among tokens except for the sink. According to the geometrical interpretation of sink, keys, and queries shown in  Figure 2c, the attention sink, which we found surprisingly almost static regardless of the input token, is just for imposing sparsity on the attention distribution.
  • It is hard to find $\mathrm{TopK}$ attention. Figure 2c also shows why searching for the Top-K keys is intrinsically costly. The keys and queries usually lie within two narrow cones with nearly opposite orientations, except for the attention sink. This significant mismatch between query and data distributions causes nearest-neighbor search methods to perform poorly.
(a) Long tailed phenomena
(b) Attention sink reshapes sparsity
(c) Geometry of attention
Figure 2: Left: Examples of long-tailed distribution in LLM. The x-axis is the fraction (or number of tokens) used in the $\mathrm{TopK}$, a.k.a. the sampling budget. Mid: Sink tokens make attention score look sparser. Right: The geometry of attention. The key of attention sink $k_{sink}$ is almost opposite to other tokens, and its orientation is surprisingly invariant with input tokens. Query states lie close to $k_{0}$, thus forming attention sink and Figure 2b. $k$ usually lies in a narrow cone that is far away from $q$. In certain heads, this geometry will result in a long-tailed distribution of attention score and difficulty searching for the $\mathrm{TopK}$ keys.
These limitations of $\mathrm{TopK}$ attention require rethinking the sparse attention approximation. Rather than only using the keys and values with the highest scores, leveraging information on the distribution can make the estimation more accurate. We approach this as a bias correction problem in sampling. Unbiased and efficient sampling has been long studied in biology [46], sociology [12] as well as machine learning [7, 13, 65], with theoretical guarantees.
Figure 3 shows that sampling values according to their corresponding attention score (we call this oracle sampling) achieves a much lower (up to 4×4$\times$4 ×) estimation error than the naive $\mathrm{TopK}$ selection. Deploying sampling estimation in attention is promising, but three challenges remain. First, how a reduction of the attention error can make a difference in downstream performance is unclear [7, 6]. Second, modeling the attention score distribution is necessary for efficient sampling, but inferring the distribution parameters requires expensive computations. Third, fully leveraging the resources of modern hardware, GPU and CPU, with a theoretically efficient algorithm is non-trivial.
@@
This paper proposes Magic samPlIng for Generation (MagicPIG), which leverages Locality sensitive hashing (LSH) sampling for efficient LLM generation. LSH is employed for sampling to approximate the attention score distribution and estimate attention output. By computing hash functions on GPU and conducting sampling on CPU, MagicPIG can allow massive hash tables and hash functions compared to prior work [37, 16], which are of vital importance for accurate estimation [6]. Following the practice of Aminabadi et al. [5], He and Zhai [27], we offload the KV cache computation, which is memory bound, to CPU to allow a larger batch or longer context. Specifically,
@@
  • In Section 4.4, we present our system design to efficiently offload attention computation on the CPU, breaking the memory limit of the GPU for serving larger batches or longer contexts. We also overcome the new challenges of computation and memory size raised by our sampling algorithm to support a larger scale of hashing tables beyond prior work [16, 37].
In Section 5, we show the empirical evaluation results of the performance of MagicPIG, demonstrating the accuracy and efficiency. While maintaining high accuracy for diverse tasks, MagicPIG can improve serving throughput by 1.93.9×1.9$\sim$ 3.9$\times$1.9 ∼ 3.9 × (A100, L40, RTX 4090) and can achieve 110ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct  [25] with 96K context. More importantly, we show that MagicPIG already outperforms $\mathrm{TopK}$ attention in the two aggregation tasks in Figure 1, suggesting that sampling indeed goes beyond $\mathrm{TopK}$ attention.
@@ 2.2 Related works
Efficient Attention. Attention approximation has been long studied. Reformer [37], KDEformer [65] and ScatterBrain [16] tackle the problem via locality sensitive hashing. These methods work in training and encoder models like BigGAN [10]. Theoretically, the error bounds and minimal workload required are continuously improved [9, 4] but have not proven to be practical for wall-clock acceleration in LLM decoding. Besides, flash-attention [22, 20, 21], flash-decoding [64, 31] and SlimAttention [28] losslessly accelerate scaled product attention operator by maximizing the utilization of hardware, which is orthogonal to our approach.
Locality sensitive hashing. Locality sensitive hashing (LSH) [7, 6] is a family of hashing functions which assigns the same hash codes for similar inputs with higher probability than others [15, 34]. LSH uses two hyper-parameters, $(K,L)$. $L$ hash tables are independently built. Each hash table has its own function $H$ which projects a high-dimension vector to an integer by concatenating $K$ random independent hash functions. In the sampling process, all vectors that share hash codes in at least one hash table with a query will be collected. SimHash [11] is the LSH family based on cosine similarity. For a vector $x\in\mathbb{R}^{d}$, SimHash generates a random hyperplane $w$ and returns $\mathrm{Sign}(w^{T}x)$. Vectors share the same sign if and only if the random projection is not in between them. For a random projection, all angles are equally likely, thus the probability that two vectors $x$, $y$ share the same sign is $p=1-\frac{\theta}{\pi}$, where $\theta=\arccos{\frac{xy^{T}}{||x||\cdot||y||}}$. If we have $L$ hash tables each with $K$ random hash functions, the probability of $y$ to be retrieved by query $x$ is $1-(1-p^{K})^{L}$.
KV Cache reduction. To get rid of memory bound introduced by KV cache thus enabling a larger batch size or serving a longer prompt, many methods are proposed to reduce the volume of KV cache. For example, H2[67], SnapKV [40] and Keyformer [2] calculate heuristics during the prefilling phase to decide which tokens to preserve for decoding phase. Quest [58] and Loki [55] do not evict KV cache but apply dynamic sparsity to reduce KV Cache loading at inference time. Besides the reduction along the dimension of sequence length, methods like KIVI [44] and QServe [41] reduce the size of KV Cache by quantization.
@@ 3 Rethinking attention sparsity
In this section, we examine $\mathrm{TopK}$ attention, which is the theoretical upper bound of prior search-based algorithms, including both static methods [67, 40] and dynamic methods [58, 55, 47]. We show that $\mathrm{TopK}$ is sub-optimal and present another attention approximation based on sampling and estimation with an oracle that improves the accuracy and/or the computation cost.
@@ 3.1 Achilles’ heel of TopK attention
As it is defined, $\mathrm{TopK}$ attention only computes the weighted average on elements with the highest attention scores. To quantify its performance, the computation budget of $\mathrm{TopK}$ attention is defined as the number of selected tokens, i.e., the $\mathrm{K}$ of $\mathrm{TopK}$. Searching-based sparse attention algorithms, like [58, 55, 62], are approximations for $\mathrm{TopK}$ attention by replacing the true $\mathrm{TopK}$ keys with the ones found by approximate searching algorithms.
However, we find significant performance degradation in downstream tasks caused by $\mathrm{TopK}$ attention as shown in Figure 1. Although $\mathrm{TopK}$ attention preserves accuracy for retrieval tasks that only require a minimal subset of the context (needle-in-a-haystack single/multikey [33]), it severely degrades for aggregation tasks that leverage the full context (common word extraction and frequent word extraction [33]). Intuitively, the information is distributed more broadly for aggregation tasks, which results in less peak attention score distribution.
@@
To better understand the attention distribution, we study the geometry of $q,k$ and make the following three observations. (1) Key states of the initial token (also known as attention sink, denoted by $k_{sink}$) remain almost the same for arbitrary input. In Figure 5a, we randomly draw $32$ samples from the vocabulary and measure the mutual cosine similarity of key states. Surprisingly, we find that the orientations of the key states of different input tokens are almost identical with a similarity $>0.99$. (2) The orientation of the center of key states (i.e. $k_{avg}=\frac{1}{n}\sum_{i=1}^{n}k_{i}$) remains stable for different input sentences. In Figure 5b, we measure the mutual cosine similarity of $k_{avg}$ of $50$ different input sentences. Although variance exists, the similarity of $k_{avg}$ is over $0.9$. (3) The orientations of $k_{avg}$ and $k_{sink}$ are almost opposite. In Figure 5c, we find that for each head, $k_{sink}$ and $k_{avg}$ has a cosine similarity between $-0.9\sim-0.8$.
These observations shape the geometry as shown in Figure 2c. The attention sink, which is static regardless of input, produces high sparsity in the attention distribution, whereas other parts are more uniformly distributed. Simply applying $\mathrm{TopK}$ will place even more weight on the sink token, thus losing contextual information. In addition, the misalignment of $q$ and $k$ also causes difficulty in search [42].
(a) $k_{sink}$
(b) $k_{avg}$
(c) Cosine similarity of $k_{sink}$ and $k_{avg}$
@@ 3.2 Estimate attention with sampling
(a) Estimation error Layer11
(b) Estimation error Layer19
(c) Downstream Comparison
@@
Existing TopK attention mechanisms ignore tokens in the KV cache with low attention scores, which introduces a bias since the ignored tokens comprise a large proportion of attention scores (Figure 2a). As a result, TopK attention achieves suboptimal performance for long-context tasks, such as information aggregation (Figure 1). Increasing the computation budget for TopK attention does help reduce the estimation error (Figure 4) since it will involve more elements in computing. However, the following question is posed:
@@
Inspired by mark and recapture [46, 48, 45, 12], we show in the following that attention output can be estimated with sampling. Using notations from Section 2.1 we can re-write attention output $o$ as the expectation of $v_{i},1\leq i\leq n$ from distribution $w$, i.e. $o=\mathbb{E}_{i\sim w}(v_{i})$, which can be estimated by the following method.
@@
This is not the lowest variance estimator but has better downstream performance (see Appendix B). We call it “oracle” because it assumes that the exact attention vector $w$ is known, which is not true for sparse attention approximations.
@@
This theorem (proved in Appendix A) theoretically guarantees a low estimation error of oracle sampling. We also present an empirical comparison between oracle sampling estimation and $\mathrm{TopK}$ attention in Figures 6a and 6b. In summary, oracle sampling estimation can reduce relative error by up to 4×4$\times$4 ×.
@@
This theorem (proved in Appendix A) shows that the computation cost of oracle sampling is usually far less than the sampling budget. In Figure 6c, we present the downstream accuracy comparison between oracle sampling estimation and $\mathrm{TopK}$ attention. The former preserves high accuracy for both tasks, even with a very small computation cost ($0.002\%$ out of 16k context, which is approximately $32$).

### v4

@@ Abstract
Large language models (LLMs) with long context windows have gained significant attention. However, the KV cache, stored to avoid re-computation, becomes a bottleneck. Various dynamic sparse or TopK-based attention approximation methods have been proposed to leverage the common insight that attention is sparse. In this paper, we first show that TopK attention itself suffers from quality degradation in certain downstream tasks because attention is not always as sparse as expected. Rather than selecting the keys and values with the highest attention scores, sampling with theoretical guarantees can provide a better estimation for attention output. To make the sampling-based approximation practical in LLM generation, we propose MagicPIG, a heterogeneous system based on Locality Sensitive Hashing (LSH). MagicPIG significantly reduces the workload of attention computation while preserving high accuracy for diverse tasks. MagicPIG stores the LSH hash tables and runs the attention computation on the CPU, which allows it to serve longer contexts and larger batch sizes with high approximation accuracy. MagicPIG can improve decoding throughput by up to 5×5$\times$5 × across various GPU hardware and achieve 54ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct model with a context of 96k tokens.
@@ 1 Introduction
Large language models (LLMs) with long context windows, such as GPT (Achiam et al., 2023), Llama (Dubey et al., 2024), and Gemini (Team et al., 2023), have gained significant attention for their ability to enhance applications like chatbots (Chiang et al., 2024), search engines (Wang et al., 2024), and video analysis (Cheng et al., 2024). However, serving long-context LLMs is highly challenging due to the unique bottleneck in auto-regressive generation—the key-value (KV) cache, which stores intermediate attention keys and values to avoid re-computation (Pope et al., 2022; Zhang et al., 2023b). Specifically, the KV cache grows linearly with both the batch size and sequence length, occupying substantial GPU memory and increasing decoding time. Moreover, the KV cache makes LLM generation extremely memory-bound, leading to underutilization of GPU computational power. For instance, an NVIDIA A100-40GB GPU can only handle a single request for Llama with a 128k context length, with nearly half of the decoding time spent accessing the KV cache, and poor GPU utilization  (He and Zhai, 2024).
Leveraging the common insight that attention is naturally sparse, dynamic sparse or $\mathrm{TopK}$-based approximation has been extensively studied (Tang et al., 2024; Singhania et al., 2024; Zhang et al., 2024; Wu et al., 2024), but three major challenges prevent a wide adoption in LLM serving systems. (1) Quality Degradation. They usually propose various strategies to approximate a subset of KV cache that yields the highest attention scores. However, $\mathrm{TopK}$ attention itself is a biased attention approximation and lacks theoretical guarantees. Figure 1 shows that even exact $\mathrm{TopK}$ attention results significantly degrade the accuracy of certain downstream tasks. (2) High Overhead. There is a large overhead to identify $\mathrm{TopK}$ attention, which becomes the bottleneck rather than the attention computation. For example, as studied in Liu et al. (2024a), naively applying a search algorithm like IVF (Douze et al., 2024) requires access over $30\%$ key states to obtain the exact $\mathrm{TopK}$, showing an unsatisfying trade-off between search accuracy and cost. (3) No Memory Saving. Although saving KV cache loading time, they cannot reduce the total memory occupied by the KV cache, which limits the maximum context and batch sizes when VRAM is scarce.
@@
  • Seemingly high sparsity is usually a consequence of an attention sink. Most of the attention scores concentrate on initial tokens (attention sink phenomenon) (Xiao et al., 2023), making the distribution look sparser. However, as shown in Figure 2(b), attention scores are distributed more uniformly among tokens except for the sink. According to the geometrical interpretation of sink, keys, and queries shown in  Figure 2(c), the attention sink, which we found surprisingly almost static regardless of the input token, is just for imposing sparsity on the attention distribution.
  • It is hard to find $\mathrm{TopK}$ attention. Figure 2(c) also shows why searching for the Top-K keys is intrinsically costly. The keys and queries usually lie within two narrow cones with nearly opposite orientations, except for the attention sink. This significant mismatch between query and data distributions causes nearest-neighbor search methods to perform poorly.
Figure 2: Left: Examples of long-tailed distribution in LLM. The x-axis is the fraction (or number of tokens) used in the $\mathrm{TopK}$, a.k.a. the sampling budget. Mid: Sink tokens make attention score look sparser. Right: The geometry of attention. The key of attention sink $k_{sink}$ is almost opposite to other tokens, and its orientation is surprisingly invariant with input tokens. Query states lie close to $k_{0}$, thus forming attention sink and Figure 2(b). $k$ usually lies in a narrow cone that is far away from $q$. In certain heads, this geometry will result in a long-tailed distribution of attention score and difficulty searching for the $\mathrm{TopK}$ keys.
These limitations of $\mathrm{TopK}$ attention require rethinking the sparse attention approximation. Rather than only using the keys and values with the highest scores, leveraging information on the distribution can make the estimation more accurate. We approach this as a bias correction problem in sampling. Unbiased and efficient sampling has been long studied in biology (Lukacs, 2009), sociology (Chen et al., 2018) as well as machine learning (Backurs et al., 2019; Chen et al., 2019; Zandieh et al., 2023), with theoretical guarantees.
Figure 3 shows that sampling values according to their corresponding attention score (we call this oracle sampling) achieves a much lower (up to 4×4$\times$4 ×) estimation error than the naive $\mathrm{TopK}$ selection. Deploying sampling estimation in attention is promising, but three challenges remain. First, how a reduction of the attention error can make a difference in downstream performance is unclear (Backurs et al., 2019, 2018). Second, modeling the attention score distribution is necessary for efficient sampling, but inferring the distribution parameters requires expensive computations. Third, fully leveraging the resources of modern hardware, GPU and CPU, with a theoretically efficient algorithm is non-trivial.
@@
This paper proposes Magic samPlIng for Generation (MagicPIG), which leverages Locality sensitive hashing (LSH) sampling for efficient LLM generation. LSH is employed for sampling to approximate the attention score distribution and estimate attention output. By computing hash functions on GPU and conducting sampling on CPU, MagicPIG can allow massive hash tables and hash functions compared to prior work (Kitaev et al., 2020; Chen et al., 2021), which are of vital importance for accurate estimation (Backurs et al., 2018). Following the practice of Aminabadi et al. (2022); He and Zhai (2024), we offload the KV cache computation, which is memory bound, to CPU to allow a larger batch or longer context. Specifically,
@@
  • In Section 4.4, we present our system design to efficiently offload attention computation on the CPU, breaking the memory limit of the GPU for serving larger batches or longer contexts. We also overcome the new challenges of computation and memory size raised by our sampling algorithm to support a larger scale of hashing tables beyond prior work (Chen et al., 2021; Kitaev et al., 2020).
In Section 5, we show the empirical evaluation results of the performance of MagicPIG, demonstrating the accuracy and efficiency. While maintaining high accuracy for diverse tasks, MagicPIG can improve serving throughput by 1.55×1.5$\sim$ 5$\times$1.5 ∼ 5 × (A100, L20, RTX 4090) and can achieve 54ms decoding latency on a single RTX 4090 for Llama-3.1-8B-Instruct  (Dubey et al., 2024) with 96K context. More importantly, we show that MagicPIG already outperforms $\mathrm{TopK}$ attention in the two aggregation tasks in Figure 1, suggesting that sampling indeed goes beyond $\mathrm{TopK}$ attention.
@@ 2.2 Related works
Efficient Attention. Attention approximation has been long studied. Reformer (Kitaev et al., 2020), KDEformer (Zandieh et al., 2023) and ScatterBrain (Chen et al., 2021) tackle the problem via locality sensitive hashing. These methods work in training and encoder models like BigGAN (Brock et al., 2019). Theoretically, the error bounds and minimal workload required are continuously improved (Brand et al., 2023; Alman and Song, 2023) but have not proven to be practical for wall-clock acceleration in LLM decoding. Besides, flash-attention (Dao et al., 2022b; Dao, 2023; Dao et al., 2022a), flash-decoding (Ye et al., 2024; Hong et al., 2024) and SlimAttention (He et al., 2024) losslessly accelerate scaled product attention operator by maximizing the utilization of hardware, which is orthogonal to our approach.
Locality sensitive hashing. Locality sensitive hashing (LSH) (Backurs et al., 2019, 2018) is a family of hashing functions which assigns the same hash codes for similar inputs with higher probability than others (Chen et al., 2020b; Jafari et al., 2021). LSH uses two hyper-parameters, $(K,L)$. $L$ hash tables are independently built. Each hash table has its own function $H$ which projects a high-dimension vector to an integer by concatenating $K$ random independent hash functions. In the sampling process, all vectors that share hash codes in at least one hash table with a query will be collected. SimHash (Charikar, 2002) is the LSH family based on cosine similarity. For a vector $x\in\mathbb{R}^{d}$, SimHash generates a random hyperplane $w$ and returns $\mathrm{Sign}(w^{T}x)$. Vectors share the same sign if and only if the random projection is not in between them. For a random projection, all angles are equally likely, thus the probability that two vectors $x$, $y$ share the same sign is $p=1-\frac{\theta}{\pi}$, where $\theta=\arccos{\frac{xy^{T}}{||x||\cdot||y||}}$. If we have $L$ hash tables each with $K$ random hash functions, the probability of $y$ to be retrieved by query $x$ is $1-(1-p^{K})^{L}$.
KV Cache reduction. To get rid of memory bound introduced by KV cache thus enabling a larger batch size or serving a longer prompt, many methods are proposed to reduce the volume of KV cache. For example, H2(Zhang et al., 2023b), SnapKV (Li et al., 2024) and Keyformer (Adnan et al., 2024) calculate heuristics during the prefilling phase to decide which tokens to preserve for decoding phase. Quest (Tang et al., 2024) and Loki (Singhania et al., 2024) do not evict KV cache but apply dynamic sparsity to reduce KV Cache loading at inference time. Besides the reduction along the dimension of sequence length, methods like KIVI (Liu et al., 2024b) and QServe (Lin et al., 2024) reduce the size of KV Cache by quantization.
@@ 3 Rethinking attention sparsity
In this section, we examine $\mathrm{TopK}$ attention, which is the theoretical upper bound of prior search-based algorithms, including both static methods (Zhang et al., 2023b; Li et al., 2024) and dynamic methods (Tang et al., 2024; Singhania et al., 2024; Mao et al., 2024). We show that $\mathrm{TopK}$ is sub-optimal and present another attention approximation based on sampling and estimation with an oracle that improves the accuracy and/or the computation cost.
@@ 3.1 Achilles’ heel of TopK attention
As it is defined, $\mathrm{TopK}$ attention only computes the weighted average on elements with the highest attention scores. To quantify its performance, the computation budget of $\mathrm{TopK}$ attention is defined as the number of selected tokens, i.e., the $\mathrm{K}$ of $\mathrm{TopK}$. Searching-based sparse attention algorithms, like (Tang et al., 2024; Singhania et al., 2024; Wu et al., 2024), are approximations for $\mathrm{TopK}$ attention by replacing the true $\mathrm{TopK}$ keys with the ones found by approximate searching algorithms.
However, we find significant performance degradation in downstream tasks caused by $\mathrm{TopK}$ attention as shown in Figure 1. Although $\mathrm{TopK}$ attention preserves accuracy for retrieval tasks that only require a minimal subset of the context (needle-in-a-haystack single/multikey (Hsieh et al., 2024)), it severely degrades for aggregation tasks that leverage the full context (common word extraction and frequent word extraction (Hsieh et al., 2024)). Intuitively, the information is distributed more broadly for aggregation tasks, which results in less peak attention score distribution.
@@
To better understand the attention distribution, we study the geometry of $q,k$ and make the following three observations. (1) Key states of the initial token (also known as attention sink, denoted by $k_{sink}$) remain almost the same for arbitrary input. In Figure 5(a), we randomly draw $32$ samples from the vocabulary and measure the mutual cosine similarity of key states. Surprisingly, we find that the orientations of the key states of different input tokens are almost identical with a similarity $>0.99$. (2) The orientation of the center of key states (i.e. $k_{avg}=\frac{1}{n}\sum_{i=1}^{n}k_{i}$) remains stable for different input sentences. In Figure 5(b), we measure the mutual cosine similarity of $k_{avg}$ of $50$ different input sentences. Although variance exists, the similarity of $k_{avg}$ is over $0.9$. (3) The orientations of $k_{avg}$ and $k_{sink}$ are almost opposite. In Figure 5(c), we find that for each head, $k_{sink}$ and $k_{avg}$ has a cosine similarity between $-0.9\sim-0.8$.
These observations shape the geometry as shown in Figure 2(c). The attention sink, which is static regardless of input, produces high sparsity in the attention distribution, whereas other parts are more uniformly distributed. Simply applying $\mathrm{TopK}$ will place even more weight on the sink token, thus losing contextual information. In addition, misaligning $q$ and $k$ also causes difficulty in search (Liu et al., 2024a).
@@ 3.2 Estimate attention with sampling
Existing TopK attention mechanisms ignore tokens in the KV cache with low attention scores, which introduces a bias since the ignored tokens comprise a large proportion of attention scores (Figure 2(a)). As a result, TopK attention achieves suboptimal performance for long-context tasks, such as information aggregation (Figure 1). Increasing the computation budget for TopK attention does help reduce the estimation error (Figure 4) since it will involve more elements in computing. However, the following question is posed:
@@
Inspired by mark and recapture (Lukacs, 2009; Owen, 2013; Lohr, 2021; Chen et al., 2018), we show in the following that attention output can be estimated with sampling. Using notations from Section 2.1 we can re-write attention output $o$ as the expectation of $v_{i},1\leq i\leq n$ from distribution $w$, i.e. $o=\mathbb{E}_{i\sim w}(v_{i})$, which can be estimated by the following method.
@@
This is not the lowest variance estimator but has better downstream performance (see Section 8). We call it “oracle” because it assumes that the exact attention vector $w$ is known, which is not true for sparse attention approximations.
@@
This theorem (proved in Section 7) theoretically guarantees a low estimation error of oracle sampling. We also present an empirical comparison between oracle sampling estimation and $\mathrm{TopK}$ attention in Figures 6(a) and 6(b). In summary, oracle sampling estimation can reduce relative error by up to 4×4$\times$4 ×.
@@
This theorem (proved in Section 7) shows that the computation cost of oracle sampling is usually far less than the sampling budget. In Figure 6(c), we present the downstream accuracy comparison between oracle sampling estimation and $\mathrm{TopK}$ attention. The former preserves high accuracy for both tasks, even with a very small computation cost ($0.002\%$ out of 16k context, which is approximately $32$). In Section 12, we provide an intuitive example to explain why sampling outperforms $\mathrm{TopK}$ in estimation.